On January ​2, 2017​, Loved Clothing Consignments purchased showroom fixtures for $11,000 ​cash, expecting the fixtures to remain in service for five years. Loved has depreciated the fixtures on a​ double-declining-balance basis, with zero residual value. On September 30,2018​, Loved sold the fixtures for $5,500 cash. Record both depreciation expense for 2018 and sale of the fixtures on September 30​,2018.

​(Record debits​ first, then credits. Select the explanation on the last line of the journal entry table. Note that 2017depreciation was recorded and posted in 2017​.)

Begin by recording Begin by recording the depreciation expense for January​ 1,2018 through September 30​, 2018.
Answer #1

Solution:

Double declining rate = (1/5)*2 = 40%

Depreciation for 2017 = $11000 *40% = $4,400

Book Value on January 1, 2018 = $11000 - $4400 = $6,600

Depreciation expense for 2018 (for 9 months upto September 30, 2018) = $6600 *40% *9/12 = $1,980

Date Accounts title Debit Credit
30-Sep-18 Depreciation expense 1980
      Accumulated depreciation 1980
(To record Depreciation for 2018)
30-Sep-18 Cash 5500
Accumulated depreciation ($4400+$1980) 6380
      Fixtures 11000
      Gain on sale of Fixtures 880